OpenStudy (anonymous):

Alicia can rown 9 mi downstream in the same time it takes her to row 5 miles upstream. She rows downtstream 2 mi/h faster than she does upstream.Find Alicias rowing rate each way.

OpenStudy (anonymous):

let her speed downstream be xmi/h and hence upstream is (x-2)mi/h time to move downstream =9/x time to move upstream =5(x-2) by the problem 9/x = 5/(x-2) or 9(x-2) =5x or 9x-18=5x or 9x-5x=18 or x=18/4 or x=4.5 mi/h for downstream speed is 4.5 mi/h and for upstream speed is 2.5 mi/h